OpenStudy (tanvirzawad):

A and B are vectors. D is del. Prove that a) D.(DxA)=0 b) Dx(DA)=0 c) D.(AxB)=B.(DxA)-A.(DxB)

OpenStudy (unklerhaukus):

a) \[\nabla \cdot(\nabla \times \vec A)=0\] b) \[\nabla \times(\nabla \vec A)=0\] c) \[\nabla\cdot(\vec A\times\vec B)=\vec B\cdot(\nabla \times \vec A)-\vec A\cdot(\nabla\times \vec B)\]

OpenStudy (anonymous):

it has a lengthy proof man...